# ClarifyCoder: Clarification-Aware Fine-Tuning for Programmatic Problem Solving

### v1

@@ 6.2. Breakdown on Clarification Categories
Model Comm. Rate Good Question Rate Pass@1 Test Pass Rate
@@
DS-Coder=DeepSeek Coder, DS-Chat=DeepSeek Chat. CodeQwen1.5=CodeQwen1.5 Chat. ClarifyCoder=ClarifyCoder w/ DeepSeek Coder as Base and loss on answers only. Test PR=Test Pass Rate, Comm.R=Communication Rate, Good Q.R=Good Question Rate. *p¡0.1; *$\leq$0.05; ***p¡0.01
@@
For the testing performance, Incompleteness receives the lowest Pass@1 (9.76% $\sim$ 46.95%) and Test Pass Rate (27.97% $\sim$ 59.36%) across all models. This supports the hypothesis that without adequate clarification, incomplete problems lead to incorrect solutions. Conversely, Inconsistency generally yields higher Pass@1 and Test Pass Rates, as models can sometimes generate correct code despite inconsistencies. For categories 1a, 1c, and 1p, most changes in Pass@1 and Test Pass Rate are statistically significant, with p-values below 0.01 for many comparisons.
@@ 6.4. Training Free Methods
In the recent years, there has been extensive studies conducted on In-Context Learning, which includes providing examples in the input prompt to the model, for it to try to make a more educated prediction. Compared to Supervised Fine-Tuning, or Parameter-Efficient Fine-Tuning, In-Context Learning does not require any additional parameter update of the model, thus, making the overall procedure less resource-intensive. Another line of research has focused on Chain-of-Thought (CoT) prompting (Wang et al., 2022), where the model is asked to “reason step-by-step”. This has proven to be an effective way to boost model performance in solving complex reasoning tasks. In order to stress the importance of Clarify-Aware Tuning, it is vital to compare its performance to that of In-Context Learning and Chain-of-Thought prompting to ensure that the increase communication capability of LLMs that we were able to achieve using Clarify-Aware Tuning would not be possible by prompting. For that reason, we have conducted the following experiments: we have constructed prompts in a 1-shot, 2-shot, 3-shot, 4-shot, and 5-shot setting, meaning that within the prompt, there are examples of questions to which the model has to answer with either clarifying questions or a direct code snippet. We also do an evaluation using Chain-of-Thought prompting in a 0-shot and 1-shot settings. These experiments were conducted using CodeLlama and DeepSeek Coder models, and the corresponding results are summarized in Figure 3
@@ 7. Discussions
Calculating loss on answer only. We also investigated the performance difference of ClarifyCoder on loss calculation during fine-tuning. We compared the clarify-aware fine-tuning where loss is calculated on both the problem and answer, and the clarify-aware fine-tuning where loss is calculated sololy on the answer. As shown in Table 3, ClarifyCoder with loss calculated only on the answer achieves the highest communication rate (63.61%) and good question rate (51.93%), outperforming the version with loss calculated on both problem and answer. This improvement in clarification metrics comes with a slight trade-off in code generation performance, with Pass@1 and Test Pass Rate decreasing by 5.92 and 6.29 percentage points respectively. However, Table 4 reveals that on the standard HumanEval benchmark, the answer-only loss version of ClarifyCoder actually performs better, with a Pass@1 of 74.85% and Test Pass Rate of 83.29%, compared to 73.01% and 82.00% for the version with loss on both problem and answer.
@@ 8. Threats of Validity
202020204040404060606060100100100100202020204040404060606060Ratio $r$ in ClarifyCoderRate (%)w/o ClarifyCoder (Comm. Rate)w/o ClarifyCoder (Good Q. Rate)w/ ClarifyCoder (Comm. Rate)w/ ClarifyCoder (Good Q. Rate)
@@ 10. Conclusion
In this research, we introduce ClarifyCoder, a novel approach to enhancing code LLMs with the ability to recognize ambiguity and request clarification. Through synthetic data generation and targeted instruction tuning, ClarifyCoder significantly improves clarification-awareness in code generation without sacrificing code generation abilities. We also developed synthetic data generation that produces clarify-aware data. Furthermore, we show that optimizing both clarify-awareness and code generation capability could be conducted in a single fine-tuning with trade-offs. This work demonstrates the importance of incorporating clarification-awareness into LLMs for improved problem-solving in real-world scenarios.
11. Acknowledgements
We would like to thank Diyi Yang for her constructive feedback and suggestions.

### v2

@@ 6.2. Breakdown on Clarification Categories
Model Comm. Rate Good Question Rate 2nd-Round Pass@1 2nd-Round Test Pass Rate
@@
[…]
@@
To compare different categories more comprehensively, we also report the pass rates in the 2nd round of HumanEvalComm evaluation. Besides the 1st round of HumanEvalComm evaluation, where Comm. Rates and Good Question Rates are calculated, the second round prompts the model to complete the code given the Q&A results (Wu and Fard, 2025). Thus, the second round of the HumanEvalComm evaluation is for code generation – the input includes: 1) coding problem in the 1st round, 2) inquiry by ClarifyCoder (same for another model of interest), 3) answer to the inquiry (provided by LLM-Judge). If there is no inquiry in the 1st round, there will be no 2nd round and the generated code in the 1st round is recorded instead. See more details in (Wu and Fard, 2025). For the testing performance, Incompleteness receives the lowest Pass@1 (9.76% $\sim$ 46.95%) and Test Pass Rate (27.97% $\sim$ 59.36%) across all models. This supports the hypothesis that without adequate clarification, incomplete problems lead to incorrect solutions. Conversely, Inconsistency generally yields higher Pass@1 and Test Pass Rates, as models can sometimes generate correct code despite inconsistencies. For categories 1a, 1c, and 1p, most changes in Pass@1 and Test Pass Rate are statistically significant, with p-values below 0.01 for many comparisons.
@@
It’s important to note that: Pass@1 (same for Test Pass Rates) in HumanEval refers to code generation ability, while Pass@1 in HumanEvalComm means the Pass@1 in postclarification (code generation in 2nd round) after the LLM has engaged in clarification dialog and received answers to its clarifying questions. In HumanEvalComm evaluation, we directly use the ClarifyCoder as the model for both the 1st round and 2nd round. Thus, the ClarifyCoder is not fine-tuned or optimized for the 2nd round input, in which the model is given the Q&A results as well as the coding task. In the second round, the same input format is used for all models to ensure fairness of evaluation. So it’s expected that the pass rates may not be optimized in the 2nd round of the model - we just want to compare the difference between categories. After investigation, we found the reason is that: ClarifyCoder is fine-tuned specifically in the clarification round (1st round - task: clarification or code generation). This is causing it to get weaker results in the postclarification round (2nd round- task: pure code generation). Future work should include a more dedicated fine-tuning for 2nd round input as well to get the pass rates for HumanEvalComm. Nevertheless, we do observe higher pass rates of ClarifyCoder in Incomplete category and ”Inconsistent + Incomplete” category. The reason could be that more useful information for incomplete problems than other categories is obtained in Q&A results for generating correct code.
@@ 6.4. Training Free Methods
Recently, there has been extensive studies conducted on in-context learning, which includes providing examples in the input prompt to the model, for it to try to make a more educated prediction. Compared to Supervised Fine-Tuning, or Parameter-Efficient Fine-Tuning, in-context learning does not require any additional parameter update of the model, thus making the overall procedure less resource-intensive. Another line of research has focused on Chain-of-Thought (CoT) prompting (Wang et al., 2022), where the model is asked to “reason step-by-step”. This has proven to be an effective way to boost model performance in solving complex reasoning tasks. In order to stress the importance of clarify-aware fine-tuning, it is vital to compare its performance to that of in-context learning and CoT prompting to ensure that the increase communication capability of LLMs that we were able to achieve using clarify-aware fine-tuning would not be possible by prompting. For that reason, we have conducted the following experiments: we have constructed prompts in a 1-shot, 2-shot, 3-shot, 4-shot, and 5-shot setting, meaning that within the prompt, there are examples of questions to which the model has to answer with either clarifying questions or a direct code snippet. We also do an evaluation using Chain-of-Thought prompting in a 0-shot and 1-shot settings. These experiments were conducted using CodeLlama and DeepSeek Coder models, and the corresponding results are summarized in Figure 4.
@@ 7. Discussions
Calculating loss on answer only. We also investigated the performance difference of ClarifyCoder on loss calculation during fine-tuning. We compared the clarify-aware fine-tuning where loss is calculated on both the problem and answer, and the clarify-aware fine-tuning where loss is calculated solely on the answer. As shown in Table 3, ClarifyCoder with loss calculated only on the answer achieves the highest communication rate (63.61%) and good question rate (51.93%), outperforming the version with loss calculated on both problem and answer. This improvement in clarification metrics comes with a slight trade-off in code generation performance, with Pass@1 and Test Pass Rate decreasing by 5.92 and 6.29 percentage points respectively. However, Table 3 reveals that on the standard HumanEval benchmark, the answer-only loss version of ClarifyCoder actually performs better, with a Pass@1 of 74.85% and Test Pass Rate of 83.29%, compared to 73.01% and 82.00% for the version with loss on both problem and answer.
@@ 8. Threats of Validity
202040406060100100202040406060Ratio $r$ in ClarifyCoderRate (%)w/o ClarifyCoder (Comm. Rate)w/o ClarifyCoder (Good Q. Rate)w/ ClarifyCoder (Comm. Rate)w/ ClarifyCoder (Good Q. Rate)
@@ 10. Conclusion
In this research, we introduce ClarifyCoder, a novel approach to fine-tuning a Code LLM to learn the clarification-seeking behavior of recognizing ambiguity when it exists and requesting clarification before generating code. Through the combination of synthetic data generation and targeted instruction tuning, ClarifyCoder significantly improves clarification-awareness of a given LLM in code generation without sacrificing code generation abilities for well-specified coding tasks that don’t need clarifications. By analyzing performance across different clarification categories and comparing with training-free baselines, we show that ClarifyCoder not only enhances clarify-awareness but also maintains efficiency and utility. Furthermore, we show that optimizing both clarity-awareness and code generation capability could be conducted in a single fine-tuning with trade-offs.